OpenStudy (jen543):

1 math question??! please help! Will fan !

OpenStudy (jen543):

Find all complex solutions of x^2+5x+7=0

OpenStudy (anonymous):

what grade

OpenStudy (jen543):

If there is more than one solution, separate them with commas

OpenStudy (jen543):

I'm in the 12th but I completely forgot how to do this, I did it a while ago

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. x2+5x+7=0 Step 1: Use quadratic formula with a=1, b=5, c=7. x= −b±√b2−4ac 2a x= −(5)±√(5)2−4(1)(7) 2(1) x= −5±√−3 2 Answer: No real solutions.

OpenStudy (jen543):

There are no solutions?

OpenStudy (jjcoolja12):

no

OpenStudy (jen543):

are you sure? because my teacher is saying there is

OpenStudy (jjcoolja12):

well its showing me there isnt